A 60.0-g object connected to a spring with a force constant of 20.0 N/m oscillates with an amplitude of 5.00 cm on a frictionless, horizontal surface.

(a) Find the total energy of the system.
mJ

(b) Find the speed of the object when its position is 1.25 cm. (Let 0 cm be the position of equilibrium.)
m/s

(c) Find the kinetic energy when its position is 3.00 cm.
mJ

(d) Find the potential energy when its position is 3.00 cm.
mJ
Answer #1

Solution,

A)

Total energy, E = 0.5 k A^2 = 0.5 x 20 x 0.05^2 = 25 mJ

B)

Speed of the object, v = w sqrt [ A^2 - x^2]

where w = angular frequency = sqrt(k/m)

x = displacement = 1.25 cm = 0.0125 m

V = 18.25 sqrt(0.05^2 - 0.0125^2) = 0.88 m/s

C)

Kietic energy, KE = E - 0.5 k x'^2

where x ' = 3 cm = 0.03 m

substitue values we get K = 16 mJ

D) Potential energy U = 0.5 k x '^2 = 0.5 x 20 x 0.03^2 = 9 mJ

Comment in case any doubt please rate my answer....